OpenStudy (anonymous):

Find all real solutions of the equation. (Enter your answers as a comma-separated list. If there is no real solution, enter NO REAL SOLUTION.) x4/3 − 14x2/3 + 48 = 0

OpenStudy (anonymous):

x^4/3-1^2/3+48=0

OpenStudy (anonymous):

do you know how to use the little equation button how you wrote it is confusing

OpenStudy (anonymous):

is it\[(x ^{4}-1)^{2}\div3+48=0\]

OpenStudy (anonymous):

\[\frac{(x ^{4}-1)^{2}}{3}+48=0\rightarrow (x ^{4}-1)^{2} = -144 \rightarrow x ^{4}-1 = \pm 12i \] I'd say NO REAL SOLUTIONS